How do I solve the linear equations using the elimination method? Can some please explain it to me?
+8x7y=12
-7x-4y=-2

Answer by maciaemi(11)

8x + 7y = 12
-7x - 4y = -2
1)First, you have to make one of the numbers the same in order to eliminate it.
So, first, multiply the whole first equation by 7 and the whole second
equation by 8. The new equation should look like this:
56x + 49y = 84
-56x - 32y = -16
2)Next, subtract the numbers. (Elimination)
49y = 84
-32y = -16
17y = 68
y = 4
3)Now that you have the answer to y, you have to find x. So now, we make the
numbers for y the same. So, multiply the first equation by 4 and the second
equation by 7.
32x + 28y = 48
-49x - 28y = -14
4)Now, subtract.
32x = 48
-49x = -14
-17x = 34
x = -2